Evaluate the integral of 5^(2y) dy


1
Expert's answer
2021-02-24T12:48:50-0500

Evaluate "\int" 52y dy

Solution

"\int" 52y dy = "\int" 25y dy

Apply exponential rule: "\int" ay dy = (ay)/(ln a) Where a = 25 in this case

Hence "\int" 25y dy = (25y)/(ln 25)

Simplyfying: "\int" 25y dy = (25y)/(2ln 5)

Adding the integral constant C: "\int" 25y dy = (25y)/(2ln 5) + C

Answer: "\int" 52y dy = (25y)/(2ln 5) + C